Grade 11Mechanics

Given a vector A = 2i + 3j and a vector B = i + j.
The component of the vector A along B is
(a) 1/(2)½ (i + j)
(b) 3/(2)½ (i + j)
(c) 5/(2)½ (i + j)
(d) 7/(2)½ (i + j)

Rajesh
9 Years agoGrade 11

7 Answers

Vikas TU
9 Years ago
The component of the vector A along B geometrically is:
Bcos(theta) = (A.B/|B|).(Bcap)
                   =5/(2)½   (i + j)
  1.  option is correct.
Pushkar Ranjan
8 Years ago
A along B is given as” A cosθ = A.B/|B|”
which formulates to give 2+3 √2 = 5/ √2
the option  (c.) is correct
 
meet kumar
7 Years ago
As we know 
ABcosø=A.B
From this eq we get
ACosø=A.B/|B|
by putting values from given data
A Cosø=5/root2
Soo C is the right option
M
7 Years ago
2square+3square whole root=5
1square+1square whole root=root 2
then multiply 
both you could get needed answer 
that is 5root2
 
Aahira
6 Years ago
The component of A along vector B=) A. B) where B^=B/b where b is the magnitude of vector B ... 
Now do A. B 
Thn find B^ wch is rdq soln
Rishi Sharma
6 Years ago
Hello students,
The problemis of vector and asked the component of A along B.
We know that,
the component of the vector A along B geometrically is :=(A.B).(Bcap)
=(A.B).(B/|B|)
=((2i +3j) . (i + j)) . ((i+j)/(2)½)
=(2+3) . ((i+j)/(2)½)
=5/(2)½ (i + j)
I hope this solution would solve your all doubts.
Thank You

LISA
5 Years ago
Magnitude of the component (let us say vecC )vector of A⃗ along B⃗ is given by,
 
|C⃗ |=A⃗ ∙B⃗ |B⃗ |=(2i^+3j^)∙(i^+j^)12+12√=2×1+3×12√=52√2 
 
Since,
 
i^∙i^=1 
 
j^∙j^=1 
 
i^∙j^=0=j^∙i^ 
 
Again, the vector along B⃗ is given by
 
C⃗ =|C⃗ |× unit vector along B⃗  
 
=|C⃗ |×B⃗ |B⃗ | 
 
=52√2(i^+j^)2√ 
 
=52i^+52j^